A student drove a distance of 135 miles at an average speed of 50 mph. How much faster would he have to drive on the return trip to save 30 minutes of driving time?
Answer by checkley77(12844)

D=RT
135=50T
T=135/50
T=2.7 HOURS.
135=R(2.7-.5)
135=2.2R
R=135/2.2
R=61.364 MPH.